Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel

P69-501

P69-501: Establishing the accuracy of asteroseismic mass, radius, and age estimates of giant stars using eclipsing binaries

P.I.: Karsten Brogaard

This is a Monitoring program using FIES.

Instrument setup

  • FIES

  • High-res fiber F4

Afternoon instructions

  • Run afternoon calibrations: fies-calibs -fib 4

Observing instructions

  • On the Queue page, make sure the OB-generated observing script is written to disk, or else 'compile'

  • Go to directory: cd ~/obs/scripts/69-501/ and do ls -lrt

  • Start the requested OB generated script


Program Tracking
Proposal 69-501
Semester P69: 20240401 - 20241001
Principal Investigator Karsten Brogaard
Programme Establishing the accuracy of asteroseismic mass, radius, and age estimates of giant stars using eclipsing binaries.
Instrument FIES
Time Allocation Multi epoch of 8 targets, ~1h/obs.
Total Nights 30.0h
Type Mon
Notes Fixed dates to be scheduled
Total Time Spent 17:40 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
69-501 2024-04-01 01:01 Mon kic9153621 FIHd010109 - FIHd010111 Target observed successfully with clear skies. We had to CTRL-C the program during acquisition multiple times because the telescope for unknown reasons moved to the wrong star when placing the target on the fiber. Hence the many entries below! Observed
69-501 2024-04-02 01:06 Mon kic4663623 FIHd020141 - FIHd020143 Clear skies, subarcsec seeing. The program had been marked as failed during the afternoon, however there was no problem observing it during the night Observed
69-501 2024-04-05 00:00 Mon - - The monitoring target was missed due to bad weather Failed
69-501 2024-04-12 01:03 Mon kic5786154 FIHd120112 - FIHd120114 There was calima, but the sky was clear. The seeing was around 2 arcsec. First try had an error during the EasyThAr part, but as it was the first thing that happened during the OB (after guiding and acquisition) we just restarted the OB and everything went smooth from there. Observed
69-501 2024-04-14 02:03 Mon epic248847494, kic4663623 FIHd140033 - FIHd140038 These were the only FIES targets of the night. Observed without difficulties. Observed
69-501 2024-04-20 00:30 Mon kic5786154 FIHd200080 - FIHd200082 Setting the observation as failed as it had to be stopped halfway because the telescope was shut down due to high humidity. Nonetheless, the data of the 30min exposure may still be worth looking at. Failed
69-501 2024-04-21 00:54 Mon epic248847494 FIHd210070 - FIHd210072 The observation was done through thin clouds, which by the end of the observation became denser, causing the guide star to be lost during the last few minutes of the exposure, ending the observation a bit sooner. Although I have set the observation as complete, please check if the data is of use. Several ctrl c's as there were several attempts to carry out the observation thwarted by the clouds. Observed
69-501 2024-04-28 00:00 Mon Failed because telescope could not be opened due to high humidity and subzero temperatures. Failed
69-501 2024-04-30 01:07 Mon kic4663623 FIHd300113 - FIHd300115 Observed with clear sky and seeing around 1-2. First attempt was accidentally observed wrong target next to correct one. Observed
69-501 2024-05-04 01:07 Mon kic5786154 FIHe040122 - FIHe040124 Observed in clear skies with seeing around 0.6" Observed
69-501 2024-05-14 01:06 Mon kic5786154 FIHe140099 - FIHe140101 Target observed with clear skies, sub-arcsec seeing. Observed
69-501 2024-05-18 01:07 Mon kic4663623 FIHe180113 - FIHe180115 Observed with clear skies and variable seeing around 1.0''. Observed
69-501 2024-05-21 01:06 Mon kic5786154 FIHe210130 - FIHe210132 Observed with thin clouds and seeing around 1 arcsec. S/N estimates from exposure meter were 48 (O9), 51 (F3), 64 (G7) and 111 (M2). Observed
69-501 2024-05-24 01:07 Mon kic4663623 FIHe240066 - FIHe240068 Observed with clear sky and better than 1 arcsec seeing. Observed
69-501 2024-05-26 01:07 Mon kic5786154 FIHe260128 - FIHe260130 The observation conditions was clear sky and good seeing. Observed
69-501 2024-05-28 00:00 Mon BD332642 ALHe280229 - ALHe280233 Shared observation from another proposal, shared here by mistake due to a typo. No time was deducted. Observed
69-501 2024-06-04 01:05 Mon kic4663623 FIHf040131 - FIHf040133 Observed with clear skies, 0.5" seeing. Spectrum might be very slightly contaminated with light from neighboring star. Observed
69-501 2024-06-07 01:06 Mon kic5786154 FIHf070100 - FIHf070102 Observed in clear skies. Observed
69-501 2024-06-08 01:05 Mon kic4663623 FIHf080075 - FIHf080077 Target observed successfully. Fiber guiding defaulted to the wrong star in the first attempt, so please ignore files FIHf080073 - FIHf080074. Observed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: March 11 2024